How much current passes through the 20-kΩ resistor when the resistors ( 20 kΩ, 30 kΩ, 60 kΩ) are connected in series with a 220 V source?
a. 2 mA
b. 11 mA
c. 20 mA
d. 110 mA
1
Expert's answer
2018-11-26T16:08:10-0500

Correct answer: a.2mA


R_Total=20+30+60=110 kΩ

I=U/R_Total =220/110000=2 mA